OpenStudy (anonymous):

If v1 = (3,7) and v2 = (-2,4), then v1*v2 is equal to which of the following? A. (-6,28) B. 34 C. (-21,-8) D. 22

OpenStudy (anonymous):

\[v_1 \times v_2 \implies (3 \times -2, 7 \times 4)\]